### v1

@@ 5.3 Control of Steps
There are three main approaches: (1) greedy selection, which generates a step and then follows it, (2) ensemble strategy, which generates a set of possible next steps, and (3) a full tree-shaped search which generates multiple options for the step, and follows them multiple steps into the future, traversing a search tree with backtracking, controlling an exponential search space. We include reinforcement learning approaches, that interactively learn an optimal policy for such a reasoning space.
5.3.1 Greedy Selection
@@
Most earlier works on prompt-based reasoning follow the greedy approach: generate a single prompt with a sequence of steps and follow them. Among the greedy reasoners are Chain-of-thought, Auto-CoT, and Zero-shot CoT. Inner Monologue and Say-Can also use greedy reasoning.
@@
In Least-to-most prompting (Zhou et al., 2022), the key idea is to break down a complex problem into simpler subproblems and then solve these in sequence, explicitly encoding them in the prompt. It is related to Complexity-based prompting. In Least-to-most, finding the answer to each subproblem is facilitated by the answers to previously solved subproblems, as in a curriculum (Bengio et al., 2009). The authors find that on symbolic manipulation, compositional generalization, and math reasoning, the Least-to-most prompting is capable of generalizing to more difficult problems than those that are given in the prompts. Figure 14 illustrates the idea.
5.3.2 Ensemble Strategy
The second kind of reasoning control is based on an ensemble of (sequences of) reasoning steps. The ensemble approach is a well-known technique in machine learning to make a strong learner out of multiple weaker learners (Sagi and Rokach, 2018; Breiman, 2001). For most problems, multiple different options for the next step exist. When all or some of these are generated and evaluated, then the best result or the consensus result can be reported as the outcome of an ensemble of steps. Various approaches have been proposed.
We already mentioned Self-consistency (Wang et al., 2022b) and Self-verification (Weng et al., 2022) in Section 5.2.1. They are popular ensemble approaches to evaluate the results of reasoning steps in prompt learning. The greedy single-path decoding used in Chain-of-thought prompting is replaced by sampling a diverse set of reasoning paths, evaluating them, and selecting the most consistent answer.
In another domain Chain-of-experts builds on Chain-of-thought with a mixture of experts ensemble for complex combinatorial operations research problems (Xiao et al., 2023b). PAL and MathPrompter also use the ensemble approach. They generate multiple steps, which are evaluated and whose answer is combined, or the best step is chosen.
The ensemble approach is a popular approach in LLM-reasoning.
5.3.3 Reinforcement Learning
In the greedy approach, a single reasoning path is generated and traversed. In reasoning, often multiple valid reasoning steps are possible, but pursuing all possibilities over multiple reasoning steps may lead to an infeasible number of possibilities.
@@ 5.3.3 Reinforcement Learning
The third kind of reasoning control is to use a full-fledged controller that can traverse a tree, or even perform reinforcement learning to do so (Sutton and Barto, 2018; Kaelbling et al., 1996; Plaat, 2022). This group of control approaches enables the most elaborate control of the reasoning process, and is used by many works, as we will see. When decomposing the problem, multiple alternative steps are generated that can be searched multiple steps into the future. Then, backtracking can be performed, allowing alternative steps to be tried.
@@
Where greedy and ensemble processes can be controlled with a prompt by the LLM, this third group is more complex, and an external algorithm is used to control the reasoning process. The external algorithms call the LLM as a subroutine prompting it to perform its tasks. This allows more complex reasoning control, but we are no longer performing prompt-based self-reasoning; control has been given to an algorithm that is external to the LLM and external to prompt-learning.
We start our discussion of control strategies with depth-first and breadth-first search, then go to beam search, and then to full reinforcement learning.
Breadth first search
A complex reasoning space can be traversed with a search algorithm. Tree-of-thoughts includes a search algorithm to dynamically follow different reasoning steps (Yao et al., 2024). When one reasoning path has been traversed, a search algorithm can backtrack, and try an alternative path. The paper describes both a breadth-first-search and a depth-first-search controller.
@@
The evaluation part in Tree-of-thoughts is performed with a prompt by the LLM. Together, the trio of generation, evaluation, and control allow systematic exploration of the space of reasoning steps with look-ahead and backtracking. The authors compare their approach to Chain-of-thought and Self-consistency. Chain-of-thought builds a reasoning out of a path of thoughts, Self-consistency creates an ensemble of thoughts, and Tree-of-thoughts constructs a tree structure. Figure 15 illustrates the different reasoning structures.444A similarly named approach is Graph-of-thoughts (Besta et al., 2024). Graph-of-thoughts allows more general reasoning graphs, providing a formal framework, where the different elements can then be specified manually.
@@
Another approach, Buffer-of-thoughts (Yang et al., 2024), goes a step further towards meta-reasoning. It introduces a meta-buffer that stores high-level thought-templates. These universal thought-templates are derived from a variety of tasks. Figure 16 compares the Buffer-of-thoughts approach to other approaches such as Chain-of-thought and Tree-of-thoughts. Buffer-of-thoughts outperforms other methods in puzzles such as Game of 24 and checkmating. Thought templates are related to metacognition (thinking about thinking), which is further discussed in Section 6.2.3.
Beam search
A related search method is Beam-search. Beam-search-for-reasoning (Xie et al., 2024) focuses on control of the space of possible reasoning paths. In some reasoning problems, this space can be very large. Beam-search solves this challenge by searching only a promising part of this space. It uses self-evaluation to control exploration and to evaluate (decode) reasoning steps. Figure 17 shows how Beam-search self-evaluation is used in multi-step reasoning.
@@
Beam search uses Program-aided-language models for math word problems (Gao et al., 2023). Using a Codex backbone (Chen et al., 2021), it surpasses the few-shot baselines by 6.34%, 9.56%, and 5.46% on the GSM8K, AQuA, and StrategyQA benchmarks, respectively.
Reinforcement learning
@@
Reinforcement learning (RL) methods are another step in the sophistication of optimization algorithms. RL learns by interactive sampling, improving its policy based on rewards from the environment (Sutton and Barto, 2018). To use reinforcement learning, the reasoning problem must be formulated as a Markov Decision Process: the agent-algorithm creates a prompt (an action), to sample a step () and get an answer (state, reward) from the environment-model (see Figure 18). The answer can then be used to improve the prompt (next action), just like reinforcement learning uses rewards to improve its policy of best actions for each state. The approaches that use reinforcement learning do so in the form of an external algorithm. No prompt has been created that performs RL by itself.
Progressive-hint-prompting (PHP) uses reinforcement learning to interactively improve prompts (Zheng et al., 2023). Figure 19 illustrates the approach.
@@
PHP is an external algorithm that calls the LLM with dynamic prompts, using previously generated answers as hints to progressively prompt the LLM toward the correct answers. It works as follows: (1) given a question (prompt), the LLM provides a base answer, and (2) by combining the question and answer, the LLM is queried and we obtain a subsequent answer. We (3) repeat operation (2) until the answer becomes stable, like a regular policy-optimizing reinforcement learning algorithm. The authors have combined PHP with Chain-of-thought and with Self-consistency. Using GPT-4, state-of-the-art performance was achieved in grade school math questions (95%), simple math word problems (91%) and algebraic question answering (79%).
Another approach that is motivated by improving answers from feedback, is Self-refine (Madaan et al., 2023). In this method, initial outputs from LLMs are improved through iterative feedback and refinement. Like PHP, the LLM generates an initial output and provides feedback for its answer, using it to refine itself, iteratively. Figures 20 and 21 illustrate the approach.
@@
Self-refine prompts the LLM in three ways: (1) for initial generation, (2) for feedback, and (3) for refinement. Note that Self-refine follows a greedy reasoning chain, learning from feedback. Self-refine has been used with GPT-3.5 and GPT-4 as base LLMs, and has been benchmarked on dialogue response generation (Askari et al., 2024), code optimization, code readability improvement, math reasoning, sentiment reversal, acronym generation, and constrained generation, showing substantial improvements over the base models.
@@
Another approach that combines reinforcement learning and LLMs is ReAct (Yao et al., 2022). Most works so far have focused on reasoning by the LLM, not on actions by an agent. A key element of reinforcement learning is that it learns a policy for an environment. The goal of ReAct is to combine progress in reasoning with action plan generation. (Or, to put it differently, most approaches use RL to improve LLM-reasoning, ReAct uses LLMs to improve RL agent policies.) ReAct uses Chain-of-thought prompt-learning as part of an RL framework that also uses external knowledge sources (Wikipedia) and finetuning, for error reduction, grounding, and for reducing hallucination. The framework allows hand-written prompts. Figure 22 shows four different prompting strategies.
On two interactive decision making benchmarks (ALFWorld and WebShop), ReAct outperforms imitation and reinforcement learning methods by an absolute success rate of 34% and 10% respectively, while being prompted with one or two in-context examples.
@@
The ReAct work has been developed further. Reflexion (Shinn et al., 2024) is built on top of ReAct. The goal is to create AI agents that learn by reflecting on failures and enhancing their results, much like humans do. Reflexion uses three language models: actor, evaluator, and reflector. It works as follows: (1) an actor generates text and actions, (2) an evaluator model scores the outputs produced by the actor, and (3) a self-reflection model generates verbal reinforcement cues to assist the actor to self-improve (see Figure 23). For the actor, Chain-of-thought (Wei et al., 2022b) and ReAct (Yao et al., 2022) can be used. Reflexion is evaluated on decision-making, reasoning, and coding tasks. Improvements of 10-20 percentage points are reported. Figure 24 shows three different prompting applications.
@@
To conclude this overview of reinforcement learning approaches, we discuss an application in the games domain. Voyager (Wang et al., 2023) is an agent for the game of Minecraft that uses an iterative prompting mechanism that generates code for embodied control. The mechanism includes Self-verification (Shinn et al., 2024). The agent has a skill library and an automatic curriculum to maximize exploration. Voyager interacts with GPT-4 through prompts. The goal of Voyager’s prompts is to discover as many diverse items in Minecraft as possible, a form of novelty search (Eysenbach et al., 2018). Voyager performs well, it shows in-context lifelong learning capability and reaches high scores by acquiring many tools (see Figure 25).
6 Discussion
@@
We have reviewed approaches for prompt-based reasoning by LLMs, highlighting techniques that have achieved a breakthrough in reasoning performance. It is time for reflection on limitations in the approaches, suggesting promising areas of future work. First we discuss issues concerning hallucination, faithful reasoning, and scaling. Then we discuss what LLMs can and cannot do. Then, we highlight connections with sequential decision processes and metacognition, and end with a research agenda.
6.1 Hallucination, Faithfulness and Scaling
Most works on reasoning in LLMs are experimental in nature. The success of in-context learning and Chain-of-thought reasoning is attracting the attention of work providing deeper insight into the reasoning processes in language models.
@@ 6.1 Hallucination, Faithfulness and Scaling
Saparov and He (2022) introduce a synthetic question/answer dataset designed to evaluate the reasoning abilities of LLMs. The work showed that LLMs are capable of reasoning to a certain degree, but that Chain-of-thought struggles with proof trees with a wide branching factor. In another study, Wang et al. (2022a) also aim to increase our understanding of how Chain-of-thought works. The authors find that it continues to work even with invalid steps in the reasoning chain. They also find that the order of the reasoning steps is important for good results. Prompts should be relevant to the question, and coherent (steps should be in the correct order). Jin et al. (2024) study the impact of reasoning step length on LLMs, finding a strong positive correlation between the length of the prompt and reasoning abilities.
These works highlight ways in which LLM-reasoning can see things that are not there. Next, we discuss works on failure modes of the Chain-of-thought approach, studying whether the reasoning of the LLM is faithful, or that it gives the right answer for the wrong reason.
6.1.1 Faithfulness

### v2

@@ 3.3. Control of Steps
The third stage is control. This stage controls how many sub-steps are generated, and how deep into the future the reasoning chain is generated. There are three main approaches: (3.3.1) greedy selection, which generates a step and then follows it, (3.3.2) ensemble strategy, which generates a set of possible next steps, and (3.3.3) a (reinforcement learning) search which generates multiple options for the steps, traversing a search tree with backtracking, controlling an exponential search space (Xu et al., 2025).
3.3.1. Greedy Selection
@@
Most earlier works on prompt-based reasoning follow the greedy approach: generate a single prompt with a single sequence of steps and follow them. Among the greedy reasoners are Chain-of-thought, Auto-CoT, and Zero-shot CoT. Inner Monologue and Say-Can also use greedy reasoning.
@@
In Least-to-most prompting (Zhou et al., 2023), the key idea is to break down a complex problem into simpler subproblems and then solve these in sequence, explicitly encoding them in the prompt, related to Complexity-based prompting. In Least-to-most the answers to previously solved subproblems help in finding the answer, as a curriculum (Bengio et al., 2009). On symbolic manipulation, compositional generalization, and math reasoning, Least-to-most prompting generalizes well, achieving 99% accuracy on a compositional generalization benchmark.
3.3.2. Ensemble Strategy
The second kind of reasoning control is based on an ensemble of (sequences of) reasoning steps. For most problems, multiple different options exist for the next step. When all or some of these are generated and evaluated, then the consensus result can be reported as the outcome of an ensemble of steps. Self-consistency (Wang et al., 2023b) and Self-verification (Weng et al., 2023) (in Section 3.2.1) are popular ensemble approaches to evaluate the results of reasoning steps, in which greedy single-path decoding used in Chain-of-thought prompting is replaced by a diverse set of paths. Taking this further, Chain-of-experts uses a mixture-of-experts ensemble for complex combinatorial problems (Xiao et al., 2023b). PAL and MathPrompter also use the ensemble approach. The ensemble approach is popular in reasoning with LLM.
3.3.3. Reinforcement Learning
@@
In reasoning, often multiple valid steps are possible, but pursuing all possibilities over multiple trajectories may lead to an infeasible number of possibilities. The third kind of reasoning control is to use a full-fledged controller that can traverse a tree, or perform reinforcement learning to do so (Sutton and Barto, 2018; Kaelbling et al., 1996; Plaat, 2022). When decomposing the problem, multiple alternative steps are generated that can be searched multiple steps into the future. Then, backtracking can be performed, allowing alternative steps to be tried.
@@
Where greedy and ensemble processes can be controlled with a prompt by the LLM, this third category is more complex, and an external algorithm is used to control the reasoning process. The external algorithms call the LLM as a subroutine prompting it to perform requested tasks. The external algorithm allows more complex reasoning control, but we are now beyond prompt-based self-reasoning: control has been given to an algorithm that is external to the LLM and external to prompt-learning.
@@
We start our discussion of control strategies with depth-first and breadth-first search, then go to beam search, and then to full reinforcement learning. A complex reasoning space can be traversed with a search algorithm. Tree-of-thoughts (Yao et al., 2024) uses breadth-first or depth-first search to dynamically follow different reasoning steps. The evaluation part in Tree-of-thoughts is performed with a prompt to the LLM. Together, the trio of generation, evaluation, and control allow systematic exploration of the space of reasoning steps with look-ahead and backtracking. The authors compare their approach to Chain-of-thought and Self-consistency on the Game of 24, Creative writing, and Mini crossword, achieving an accuracy of 74% on a Game of 24 task. The other tasks are evaluated qualitatively. Figure 5 illustrates the different reasoning structures.333A similarly named approach is Graph-of-thoughts (Besta et al., 2024). Graph-of-thoughts allows more general reasoning graphs, providing a formal framework, where the different elements can then be specified manually.
@@
Another approach, Buffer-of-thoughts (Yang et al., 2024), goes a step towards metareasoning (Gao et al., 2024). It introduces a meta-buffer that stores high-level thought-templates. These thought-templates are derived from a variety of tasks. Figure 5 compares the Buffer-of-thoughts approach to other approaches such as Chain-of-thought and Tree-of-thoughts. Buffer-of-thoughts outperforms other methods in puzzles such as Game of 24 (by 11%) and checkmating (by 51%). Thought templates are related to metacognition (thinking about thinking), which is further discussed in Section 4.3.3.
@@
A related search method is Beam-search-for-reasoning (Xie et al., 2024). When the space of possible reasoning paths is large, Beam-search searches a promising part of this space. It uses self-evaluation to control exploration and to evaluate (decode) reasoning steps. Beam search uses Program-aided-language models for math word problems (Gao et al., 2023). Using a Codex backbone (Chen et al., 2021), it surpasses the few-shot baselines by 6.34%, 9.56%, and 5.46% on the GSM8K, AQuA, and StrategyQA benchmarks, respectively.
@@
Reinforcement learning is another step in the sophistication of optimization algorithms. It learns by interactive sampling, improving its policy based on rewards from the environment (Sutton and Barto, 2018). To use reinforcement learning, the reasoning problem is formulated as a Markov Decision Process: the agent-algorithm creates a prompt (an action), to sample a step () and get an answer (state, reward) from the environment-model. The answer can then be used to improve the prompt (next action), using the rewards to improve its policy of best actions for each state. The approaches that use reinforcement learning also do so in the form of an external algorithm.
@@
Progressive-hint-prompting (PHP) uses reinforcement learning to interactively improve prompts (Zheng et al., 2023). PHP calls the LLM with dynamic prompts, using previously generated answers as hints, to progressively prompt the LLM towards the correct answers. It works as follows: (1) given a question (prompt), the LLM provides a base answer, and (2) by combining the question and answer, the LLM is queried and obtains a subsequent answer. We (3) repeat operation (2) until the answer becomes stable, as a regular policy-optimizing reinforcement learning algorithm. The authors have combined PHP with Chain-of-thought and with Self-consistency. Using GPT-4, state-of-the-art performance was achieved on grade school math questions (95%), simple math word problems (91%) and algebraic question answering (79%).
@@
Another approach that is motivated by improving answers from feedback, is Self-refine (Madaan et al., 2023). Like PHP, the LLM generates an initial output and provides feedback for its answer, using the LLM to refine itself, iteratively. Self-refine prompts the LLM in three ways: (1) for initial generation, (2) for feedback, and (3) for refinement, following a greedy reasoning chain. Self-refine has been used with GPT-3.5 and GPT-4 as base LLMs, and has been benchmarked beyond math word problems on dialogue response generation (Askari et al., 2024), code optimization, code readability improvement, math reasoning, sentiment reversal, acronym generation, and constrained generation, showing substantial improvements over the base models (typically around 30 percentage points, see Table 2).
@@
Another approach that combines reinforcement learning and LLMs is ReAct (Yao et al., 2023). Most works focus on reasoning by the LLM, not on actions by an agent. The goal of ReAct is to combine progress in reasoning with action plan generation. (Or, to put it differently, other approaches use RL to improve LLM-reasoning, ReAct uses LLMs to improve RL agent policies.) ReAct uses Chain-of-thought prompt-learning as part of an RL framework that also uses external knowledge sources (Wikipedia) and finetuning; for error reduction, grounding, and for reducing hallucination. The framework allows hand-written prompts. On two interactive decision making benchmarks (Alfworld and WebShop), ReAct outperforms imitation and reinforcement learning methods by an absolute success rate of 34% and 10% respectively.
@@
The ReAct work has been developed further. Reflexion (Shinn et al., 2024) creates AI agents that learn by reflecting on failures and enhance their results, much like humans do. Reflexion uses three language models: actor, evaluator, and reflector. It works as follows: (1) an actor generates text and actions, (2) an evaluator model scores the outputs produced by the actor, and (3) a self-reflection model generates verbal reinforcement cues to assist the actor to self-improve (see Figure 6). For the actor, Chain-of-thought and ReAct can be used. Reflexion is evaluated on decision-making, reasoning, and coding tasks. Improvements of 10-20 percentage points are reported.
The reinforcement learning approaches that we discussed so far—React, Self-refine, Tree-of-thoughts, Buffer-of-thoughts, Reflexion—use external algorithms to manage state for the prompt improvement loop. A more elegant solution would be to perform reinforcement learning fully in-context, within-prompt. Indeed, Krishnamurthy et al. (2024) explicitly asked the question whether LLMs can explore in context. This is the goal of the Algorithm-of-thoughts (AoT) approach (Sel et al., 2023). Following work by Lee et al. (2023); Wang et al. (2024c); Demircan et al. (2024) that showed that transformer architectures can be pretrained and finetuned to perform in-context reinforcement learning, AoT aims to do so in general LLMs such as GPT-4, Claude and Gemini 1.5. They achieve results comparable to Tree-of thoughts on GSM8K, StrategyQA, and Crosswords. Achieving these results with in-context RL is a promising result for in-context learning. Other search-like in-context algorithms are studied by Schultz et al. (2024) and Kempinski et al. (2025) (Game-of-thoughts). Further work in games finds that LLMs struggle with the knowing-doing gap and the difference between using and mentioning game concepts (Paglieri et al., 2024; van Wetten et al., 2025; Cloos et al., 2024).
@@
To conclude this overview of reinforcement learning approaches, we discuss an application in the games domain. Voyager (Wang et al., 2024a) is an agent for the game of Minecraft that uses an iterative prompting mechanism that generates code for embodied control. The agent includes self-verifcation an a skill library to maximize exploration. The goal of is to discover diverse items in Minecraft, a form of novelty search (Eysenbach et al., 2019). Voyager performs well, reaching high scores by acquiring many tools (see Figure 7) 15 times faster than the baseline.
The applications of reinforcement learning in LLM reasoning are many, and the connections run deep (Pternea et al., 2024). Wang et al. (2024b) use the similarity between RL timesteps and LLM reasoning steps to jointly train a value function together with the LLM policy by optimizing the soft Bellman equation, achieving 85% accuracy on GSM8K and 81% on Alfworld. Guo et al. (2025); Du et al. (2025) replace supervised finetuning by reinforcement learning, integrating it with reasoning, achieving large efficiency gains in the training pipeline in the DeepSeek r1 and Kimi models.
4. Discussion
@@
We have reviewed many reasoning approaches. It is now time to reflect on the approaches, to discuss limitations, and to look for promising areas of future work. First we discuss how to choose an approach for an application. Next, we discuss issues concerning hallucination, faithful reasoning, and scaling. Then we discuss what LLMs can and cannot do. Then, we highlight connections with sequential decision processes, metacognition, and programming languages, and we end with a research agenda.
4.1. Matching Methods and Applications
This survey has discussed different approaches, applications, and benchmarks. It is often difficult to directly compare benchmark results between individual papers, due to differences in measurement setup. For example, Kamoi et al. (2024) critically compare self-reflection studies. We can provide, however, general guidance on what approach to use for different types of applications.
For simple reasoning situations, that require a single sequence of reasoning steps, Chain-of-thought is typically used. When the error rate is high, verification methods such as Self-consistency are popular. Indeed, this combination is used in many modern reasoning LLMs.
For combinatorial puzzles and games that would traditionally require backtracking, tree-search or reinforcement learning, solution methods need to be able to keep track of the enumeration state. However, an LLM without an external algorithm often fails to keep track of this state correctly between calls to the model. An approach such as Tree-of-thoughts, or ReAct, should be considered. Here, the state is managed externally, and the LLM is called for application-dependent functional processing. Various in-context reinforcement learning methods are being developed where state is managed by the LLM, as in, for example, Algorithm-of-thoughts. When performance is not sufficient, finetuning or pretraining for the domain at hand can be used.
For complex problems, where a sequence of steps or external control fail to perform well, other reasoning approaches are needed. We note that LLMs are good at generating computer code for many known problems (Chen et al., 2021). Hence, when the preceding approaches do not work, the LLM can be asked to generate a program for the problem. In this approach, the LLM acts as a coding agent for its user. The LLM generates a program or a sequence of commands, which are then executed by an external tool, such as an interpreter, a robot, a planner, or a logic solver. In robotics, vision-language-action models have achieved impressive results (Kim et al., 2024), but also strong results in logistics (Bohnet et al., 2024), finance and medicine are reported (Plaat et al., 2025).
4.2. Hallucination, Faithfulness and Scaling
@@
Reasoning by LLMs is not error-free, and many studies aim to provide deeper insight into the reasoning processes in language models. Saparov and He (2023) introduce a synthetic question/answer dataset designed to evaluate the reasoning abilities of LLMs. The work showed that LLMs are capable of reasoning to a certain degree, but that Chain-of-thought struggles with proof trees with a wide branching factor. In another study, Wang et al. (2023a) aim to increase our understanding of how Chain-of-thought works. The authors find that the order of the reasoning steps is important. Prompts should be relevant to the question, and coherent (steps should be in the correct order). Jin et al. (2024) also study the impact of reasoning step length on LLMs, again finding a strong positive correlation between the length of the prompt and reasoning abilities. Next, we discuss works on errors in the Chain-of-thought approach, studying whether the reasoning of the LLM is faithful, or that it gives the right answer for the wrong reason.
4.2.1. Faithfulness